OpenStudy (anonymous):

In which interval does f(x) = -2x^2 - 3x - 4 have a maximum value?

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

its max is at x = -3/4

OpenStudy (anonymous):

It is a quadratic with a negative parabolic curve which continues to infinity in both directions. Do you mean what is the maximum value of the range?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its result is (-1, 0) but i don't know why?

OpenStudy (anonymous):

ok... Differentiate the function giving -4x -3 [for maxima or minima, in this case maxima] so x = -3/4

OpenStudy (anonymous):

but how we know max = (-1 , 0)?

OpenStudy (anonymous):

because that is a range of possible x's from -1 to 0. The value for x = -3/4 lies in that range

OpenStudy (anonymous):

basically, x = -3/4 is between -1 and 0

OpenStudy (anonymous):

ohhh ok Thanks